OpenStudy (anonymous):

Are these lines parallel perpendicular or neither? Y=-4+3 -2x+8y=5

OpenStudy (ajanijones):

Noo I im pretty sure

OpenStudy (ajanijones):

wait

OpenStudy (anonymous):

Y=-4+3 where is the x

OpenStudy (anonymous):

y = -3/4x - 2 6x + 8y = -5 Slope Intercept form of the equation for a straight line is y = mx + b where m is the slope and b is the y intercept. First equation is already in that form. In the second equation, solve for y in the equation of: 6x + 8y = - 5 Subtract 6x from both sides of this equation to get: 8y = -6x - 5 Divide both sides of this equation to get: y = (-6x - 5)/8 which becomes: y = (-6/8)*x - (5/8) Slope of the first line is -(3/4) Slope of the second line is -(6/8) -(3/4) is the same as -(6/8) so the lines are parallel.